# Multimodal Banking Dataset: Understanding Client Needs through Event Sequences

### v1

@@
Abstract
Financial organizations collect a huge amount of data about clients that typically has a temporal (sequential) structure and is collected from various sources (modalities). Due to privacy issues, there are no large-scale open-source multimodal datasets of event sequences, which significantly limits the research in this area. In this paper, we present the industrial-scale publicly available multimodal banking dataset, MBD, that contains more than 1.5M corporate clients with several modalities: 950M bank transactions, 1B geo position events, 5M embeddings of dialogues with technical support and monthly aggregated purchases of four bank’s products. All entries are properly anonymized from real proprietary bank data. Using this dataset, we introduce a novel benchmark with two business tasks: campaigning (purchase prediction in the next month) and matching of clients. We provide numerical results that demonstrate the superiority of our multi-modal baselines over single-modal techniques for each task. As a result, the proposed dataset can open new perspectives and facilitate the future development of practically important large-scale multimodal algorithms for event sequences.
@@ Abstract
1 Introduction
The ever-growing digital footprint of consumers has led to an explosion of data collected by businesses across all industries. Financial institutions, in particular, gather a vast amount of information about their clients over time, encompassing various aspects of their financial activities. This data, spanning extended periods, is typically annotated with temporal information, forming what is known as event sequences [8, 14, 23, 37]. Description of an event usually contains several heterogeneous fields, numerical and categorical, and processing of event sequence is a separate problem in machine learning different from techniques applied to time series or texts.
An important property of event sequences is that these data are often gathered from multiple sources or channels, rendering multimodal. The comprehensive analysis of such data in a joint way is critical for informed decision-making processes within enterprises. The scarcity of publicly available datasets in this domain hampers research and development efforts. Indeed, the significant progress in multimodal data for natural language processing, computer vision, audio, etc., is due to the appearance of large-scale datasets, such as LAION, DataComp, etc. [16]. Unfortunately, though several datasets of event sequences are used in research, e.g., credit card transactions [30] or MIMIC [21], they are either small or contain only one modality. Thus, a gap exists in methodologies designed to tackle the complexity of multimodal event sequence data.
@@ 1 Introduction
To bridge this gap, this paper introduces the Multimodal Banking Dataset (MBD), an unprecedented open-source resource encompassing extensive multichannel event sequence data of banking corporate clients. This dataset is the largest of its kind, featuring detailed records of approximately 1.5 million clients across four distinct modalities: money transfers (about 950 million events), geo position data (around 1 billion events), technical support dialog embeddings (approximately 5 million entries), and monthly aggregated bank product purchases categorized into four types. Each modality encompasses roughly one or two year of historical, time-annotated data, making it a rich resource for analyzing the dynamics of client behavior over time.
@@
The MBD dataset enables the exploration of several critical business problems in a multimodal context. Specifically, it facilitates research in such tasks as purchase prediction for the following month (campaigning), and matching of different modalities for the same clients. These problems can all be addressed by leveraging the temporal and spatial attributes inherent in the dataset. Moreover, we provide benchmarks for these tasks using the MBD. Importantly, all the data are anonymized to protect client privacy, and our research confirms that the anonymization process preserves the consistency of model performance between original and anonymized data. This paper underscores the potential of MBD as a foundational tool for advancing research in multimodal sequential data analysis, fostering innovation, and enhancing decision-making processes in the financial sector.
2 Dataset
@@
Modern innovative banking institutions actively develop AI technologies for customizing their human-oriented technologies and making everyday decisions. A superior level of technologies will lead to new cases of customers’ experience, which should form a competitive advantage of services provided, including the speed, accuracy, and price of customer services, including personal credit conditions, individual finance strategy, etc.
@@
One of the main benefits of using AI is the ability to analyze large amounts of customer data. This helps banks better understand their customers’ needs and offer them the most suitable products and services. Additionally, AI can protect customers from fraud and prevent financial losses. More accurate forecasting of financial risks associated with lending or investing allows us to provide more favorable conditions to more reliable clients.
This strategy is more effective if there is a lot of data, so banks strive to accumulate as much information as possible.
@@
Being a team of a bank that stores petabytes of data about bank processes and clients, we understand the urgent needs of fintech data scientists in large sets of publicly available temporal data from various sources (bank transactions, client locations, purchased products, etc.) to drive innovation and scientific discovery. Unfortunately, the number of appropriate datasets is limited because providing such data comes with certain risks. Typically, banks are wary of sharing their data due to potential leaks of confidential information or violations of data protection regulations. In addition, the data is considered to have commercial value, and companies do not want to disclose it. To mitigate these risks, it is required to take all necessary steps to ensure data security and remove any identifying information. This allows information to be shared without violating client confidentiality, but this procedure requires significant effort from engineers, managers, and lawyers. Though there exist a small number of properly anonymized banking datasets, such as credit card transactions [30] or AlphaBattle [2], to the best of our knowledge, there are no publicly available large multimodal temporal datasets for banks.
@@
Thus, in this paper, we introduced the first large-scale multimodal banking dataset to support future research on multimodal techniques for event sequences. In particular, we selected several practically important tasks, such as campaigning, i.e., prediction if a client would purchase some of four rather popular products in the next month. Each client is described by sequences of typical bank data: transactions, geo positions where the customer used the bank application, and dialogues with technical support. These data sources ideally highlight the main difficulties in developing multimodal models, namely, asynchronous events in different modalities, various intensities of events, rare/irregular events, and even the absence of some modalities for many clients. Based on our dataset, the researchers will be able to fully take into account cross-modal connections of sequences from multiple sources at the level of individual events.
@@
The dataset was collected using the following procedure. At first, we selected a complete sample of clients for two year (2021, 2022) to be able to cover all seasons. Among all customers who had the opportunity to purchase at least one of four products during 2022, we randomly selected 2,186,230 clients, among which 1M customers are labeled by monthly aggregated purchases of each of four products in each month. For these clients, we collected 947,899,612 bank transactions, 1,117,213,760 geo position events, and 5,080,781 dialogues with technical support. Our data raise typical practical challenges for training multimodal models. For example, many clients do not have all three modalities simultaneously because they can never make a transaction, call tech support, or leave their geo trace while running the bank application. Next, all the data were properly anonymized to guarantee the confidentiality and privacy of customer information. As a result, it becomes impossible to recover back to real clients from our anonymous data. We will show in the experimental study that such anonymization still let us extract valuable information about clients.
@@
To demonstrate the complete transformation of data from various sources, Fig. 1 briefly shows example data for each modality, and the temporal structure of the data for a campaigning model is presented. Let us introduce further details for each modality in the following subsections.
2.1 Modalities
@@
1. Bank transactional data are financial operations (events) carried out between different clients. The sequence of transactions can uniquely characterize the client [8], so this modality plays one of the most important roles in financial planning and recommendations. Thus, the main component of our MBD dataset is each client’s transactional history, represented by an event with a timestamp and various attributes of the anonymized counterparty. Clients have 638 transactions on average. Fig. 2 in Appendix A contains the detailed distribution of clients for different lengths of the transaction history. The list of attributes is shown in Table 4.
@@
2. Dialogues. The dialogue data are transcriptions of customer calls to technical support and negotiations between the client and its manager. It is an extremely important source of information about client needs and problems [11]. The audio utterance is fed into a commercial Speech-to-Text algorithm. Personal information, e.g., the client’s name, is detected in the text and masked. To further anonymize the dialogue, we feed its text into a pre-trained NLP model111 https://huggingface.co/ai-forever/ruBert-base and save the resulting embeddings of size 768 in dialogue modality. An example of dialogue data is presented in Fig. 3 in Appendix A. Only 46% of customers contact support and have records of conversations, 98% of them have no more than 10 dialogues per year. Fig. 4 contains the detailed distribution of clients for various numbers of dialogues.
@@
3. Geostream data contains a sequence of geo-coordinates of a client. To anonymize this modality, the coordinates are encoded using geohashes222https://pypi.org/project/pygeohash/, a geocoding system that converts a geographic location into a short string. Each unique geohash corresponds to a region on the Earth’s surface. It is possible to adjust the accuracy and size by removing characters from the end of the code. In our dataset, the coordinates were encoded with a precision of 4, 5, and 6 characters, representing cells of different sizes on the map. In our dataset, there are 43,999 distinct values of geohash_4, 347,698 numbers of geohash_5, and 2,264,404 most precise locations (geohash_6).
An example of geo data is presented in Fig. 5 in Appendix A. The clients have an average of 910 geo-points in the time period considered. Detailed information about the distribution of their length over the clients is available in Fig. 6.
@@
4. Products purchases. Our dataset is a valuable resource for analyzing the bank customers’ needs. In particular, it contains information about purchases of four different banking products for each month of 2022. To predict a purchase in a certain month, it is necessary to take events (transactions, geo, dialogues) strictly before the beginning of this month. Therefore, the date range for the purchases dataset is shifted by 1 month, i.e., information is available from February 1, 2022, to January 31, 2023. The campaigning task is a multi-label classification problem, i.e., we store a binary label for each product that indicates whether it was purchased by a customer in a certain month. Example of data about purchases is shown in Fig. 8 (Appendix A). The peculiarity of this dataset is its imbalance, which is specific to this type of business task: 81% of clients have no purchases, 15% have one, and the remaining 4% have two or more purchases. A historical overview over 12 months allows us to model the customer behavior dynamic and predict the date of purchase more accurately.
2.2 Data anonymization
@@
Our dataset contains no personal or confidential information whatsoever. Nevertheless, the event sequences are detailed enough that it could be possible to compare individuals from the publicly accessible portion of the dataset with the original proprietary data. To mitigate this risk, noise has been introduced to the data, ensuring that such comparisons and identification are impossible. The noise patterns were selected by our bank’s internal security department. These patterns are applied locally, preserving the overall structure of the data. The specific parameters of the noise are not disclosed to prevent potential attacks on the dataset.
@@
All ID fields were hashed with a random salt. All categorical field values was mapped to enumerated indexes. Random noise was added to numerical fields. The noise was added to date fields with preserving hour of original date, which may be the cause the shuffle of local sequence. The dialogue embedding space was divided into regions, which were then shuffled.
3 Benchmark
In this paper, we formulate two practically important tasks using our new dataset: campaigning and client matching. This section contains details about baseline methods and testing protocols.
3.1 Methods
To establish performance baselines, we implement several widely-adopted architectures. Our approach prioritizes unsupervised and semi-supervised methods [10], enabling the training of a general-purpose encoder on unlabeled sequential data. Additionally, we incorporate supervised methods that allow for the immediate training of the encoder in a fully supervised manner.
The following techniques are implemented in our benchmark to extract features for Transactions and Geostream modalities:
1.
Aggregation Baseline that contains hand-crafted aggregation statistics [8]: 3,000 and 4 features in total for sequences of transactions and geodata, respectively.
2.
CoLES (Contrastive Learning for Event Sequences), a self-supervised contrastive model [8] specially developed to obtain representations of such event sequences as bank transactions. The sequence encoder is a GRU (Gated Recurrent Unit) with a hidden size 256.
3.
Two Tabular Transformers from IBM [30]. The first model, TabBERT, adapts BERT to event sequences such as bank transactions. The second model, TabGPT, was initially proposed to generate synthetic tabular sequences. Both models extract 256-dimensional embeddings of an input event sequence. After, we pool output embeddings of the client in result embedding of size 1024, calculating min, max, mean, and std.
To obtain representation of a sequence of dialogues, we implemented several straightforward techniques to aggregate the sequence of embeddings of each dialog of a client: 1) mean pooling of all embeddings; and 2) use only the most recent embedding for the date of interest.
To demonstrate the potential of multimodal processing of event sequences, we implemented traditional fusion techniques, namely, 1) blending, i.e., a weighted sum of scores (estimates of class posterior probabilities) from single-modal classifier, and 2) late fusion, i.e., learning a classifier for concatenated embeddings from all modalities [20]
For supervised methods, we use the following architectures: for unimodal data, we utilize an GRU encoder, while for multimodal data, we employ different GRU encoders for each modality and embeddings from all encoders are concatenated. The resulting neural network is trained using binary cross-entropy (BCE) loss in multilabel setting. For the GRU encoders we used hidden size 32.
3.2 Tasks
3.2.1 Campaigning task

### v2

@@
Abstract.
Financial organizations collect a huge amount of temporal (sequential) data about clients, which is typically collected from multiple sources (modalities). Despite the urgent practical need, developing deep learning techniques suitable to handle such data is limited by the absence of large open-source multi-source real-world datasets of event sequences. To fill this gap, which is mainly caused by security reasons, we present the first industrial-scale publicly available multimodal banking dataset, MBD, that contains information on more than 2M corporate clients of a large bank. Clients are represented by several data sources: 950M bank transactions, 1B geo position events, 5M embeddings of dialogues with technical support, and monthly aggregated purchases of four bank products. All entries are properly anonymized from real proprietary bank data, and the experiments confirm that our anonymization still saves all significant information for introduced downstream tasks. Moreover, we introduce a novel multimodal benchmark suggesting several important practical tasks, such as future purchase prediction and modality matching. The benchmark incorporates our MBD and two public financial datasets. We provide numerical results for the state-of-the-art event sequence modeling techniques including large language models and demonstrate the superiority of fusion baselines over single-modal techniques for each task. Thus, MBD provides a valuable resource for future research in financial applications of multimodal event sequence analysis.
@@ Abstract.
1. Introduction
In the key tasks in the banking industry, such as campaigning, fraud detection, credit risk assessment, customer segmentation, and personalized recommendations, the performance heavily depends on the processing of the client’s financial and non-financial activities, e.g., transactions, dialogues with sales managers, product purchase history, etc. (Babaev et al., 2022; Rosset et al., 2001). This data, spanning extended periods, is typically annotated with temporal information, forming what is known as event sequences (Udovichenko et al., 2024; Yeshchenko and Mendling, 2022; Kolosnjaji et al., 2016; Weiss and Hirsh, 1998; Guo et al., 2020). An event is described by several heterogeneous fields, numerical and categorical. An essential property of event sequences is that these data are often gathered from multiple sources or channels, rendering them multimodal.
The success of financial organizations strongly depends on their availability to deeply analyze such multi-source heterogeneous event sequences for building classification and predictive models. However, existing multimodal models (Xu et al., 2023; Zhang and Yan, 2023) cannot be directly applied to such event/tabular data due to their significant difference from audio, images, texts, and regular time series. Unfortunately, despite the urgent business needs, the progress in the development of multimodal techniques for multi-source event sequences is limited by the absence of large-scale datasets. Indeed, though several datasets of event sequences are used in research, e.g., credit card transactions (Padhi et al., 2021) or MIMIC (Johnson et al., 2023), they are either small or contain only one modality. Thus, tackling the complexity of multimodal event sequence data is still very challenging.
@@ 1. Introduction
To bridge this gap caused mainly by security reasons in the fintech industry, this paper introduces the Multimodal Banking Dataset (MBD), an unprecedented open-source resource encompassing extensive multichannel event sequence data of banking corporate clients. It is the largest dataset of its kind, featuring detailed records of approximately 2 million clients across four distinct modalities: money transfers (about 950 million events), geo position data (around 1 billion events), technical support dialog embeddings (approximately 5 million entries), and monthly aggregated bank product purchases categorized into four types. Each modality encompasses roughly one or two years of historical, time-annotated data, making it a rich resource for analyzing the dynamics of client behavior over time.
@@
The MBD dataset enables the research of several critical business problems in a multimodal context, such as future purchase prediction (campaigning) and matching different modalities for the same clients. To support these tasks, we provide benchmarks using the MBD and other existing financial datasets of much smaller size. To ensure client privacy, all data in MBD has been rigorously anonymized, allowing researchers to work with the dataset without compromising confidentiality. Moreover, our experiments confirm that our anonymization procedure preserves the consistency of model performance between original and anonymized data.
2. Related Works
2.1. Financial data
Banks’ multitude of services and processes generates a variety of data that can be considered as modalities. Early works (Moro et al., 2014; Mancisidor et al., 2021) use feature processing techniques that remove multimodal complexity and present data in tabular form. A wide range of financial aggregates are presented in the American Express dataset (ame, [n. d.]) as a sequence of historical slices. The development of deep learning methods has led to the ability to work with complexly structured data, such as sequences of events. Quite a few datasets (age, [n. d.]; ros, [n. d.]; alp, [n. d.]), mostly unimodal, presented mainly at ML competitions. To work with such data, both supervised (Ala’raj et al., 2022; Babaev et al., 2019; Wang and Xiao, 2022) and unsupervised methods (Padhi et al., 2021; Bazarova et al., 2025; Skalski et al., 2023) are used. A multimodal financial sequential dataset was introduced in DataFusion 2022 competition (dat, [n. d.]). There are two sequential modalities, transaction and web clickstream, and two downstream tasks: matching and education level prediction. However, this is an extremely small dataset of 22K clients, and no accurate baseline model is available.
2.2. Other event sequence domains
Temporal point process (Mei and Eisner, 2017; Zhuzhel et al., 2023) model streams of discrete events in continuous time by constructing a neurally multivariate point process. The authors use a large collection of single-modal datasets from different types of modalities: Media (Retweets (Zhao et al., 2015), MemeTrack (Leskovec and Krevl, 2014), Amazon (ama, 2018), IPTV (Luo et al., 2014)), Medical (MIMIC-II (Johnson et al., 2016)), Social (Stack Overflow (Leskovec and Krevl, 2014), Linkedin (Xu and Zha, 2017)) and Financial (Transaction (Fursov et al., 2021)) data. Multimodal medical record datasets, MIMIC-IV (Johnson et al., 2023) is used in EventStreamGPT (McDermott et al., 2024). The authors propose a GPT-like approach for continuous-time event sequences. This dataset’s structure is close to our MDB. However, unlike medical information, financial data contains longer chains of events and more regular patterns, and individual transactions are less informative.
2.3. Geostream and dialogues
Geodata is used for various tasks. One of the uses of geo is a visualization of analytics on a map (Hao et al., 2011). However, geo is used here not as a separate modality but as additional tags to the mainstream of tweets. Mobile marketers (Baye et al., 2024) use geo-targeting for pricing and send personalized recommendations. In (Verma et al., 2020), geo hashes are used for user mobility detection and prediction.
We encode our dialogue entries via a pretrained NLP model. Such models have already been used (Hassan et al., 2019) for text anonymization tasks. Embeddings preserve the meaning of the text, which was shown in (Vaswani et al., 2017). Pre-trained text embeddings can capture text sentiment and improve text-to-speech models (Hayashi et al., 2019).
3. Proposed Dataset
3.1. Overview of the MBD dataset
Pipeline for processing original data sources for solving campaigning task.
| Dataset | # Clients | Downstream Tasks | # Events | Class Balance | Modalities | ||||
| Datafusion (dat, [n. d.]) | 22K |
146M | Imbalanced | Transactions, Clickstream | ||||
| Alphabattle (alp, [n. d.]) | 1.5M | Binary classification | 443M | Imbalanced | Transactions | ||||
| Age (age, [n. d.]) | 50K | Multiclass classification | 44M | Balanced | Transactions | ||||
| Rosbank (ros, [n. d.]) | 10K | Binary classification | 1M | Imbalanced | Transactions | ||||
| Credit Card Transaction (Padhi et al., 2021) | 2K |
2M | Highly Imbalanced | Transactions | ||||
| MBD-small (ours) | 70K |
80M | Imbalanced |
| MBD (ours) | 2M |
2B | Highly Imbalanced |
@@ 3.1. Overview of the MBD dataset
Modern innovative banking institutions actively develop AI technologies to customize their human-oriented technologies and make everyday decisions. A superior level of technology will lead to new customer experience cases, which should form a competitive advantage for the services provided, including the speed, accuracy, and price of customer services, including personal credit conditions, individual finance strategy, etc.
@@
One of the main benefits of using AI is the ability to analyze large amounts of customer data. This helps banks better understand their customers’ needs and offer them the most suitable products and services. Additionally, AI can protect customers from fraud and prevent financial losses. More accurate forecasting of financial risks associated with lending or investing allows us to provide more favorable conditions to more reliable clients. This strategy is more effective if there is a lot of data, so banks strive to accumulate as much information as possible.
@@
Being a team of a bank that stores petabytes of data about bank processes and clients, we understand the urgent needs of fintech data scientists in large sets of publicly available temporal data from various sources (bank transactions, client locations, purchased products, etc.) to drive innovation and scientific discovery. Unfortunately, the number of appropriate datasets is limited because providing such data comes with certain risks. Typically, banks are wary of sharing their data due to potential leaks of confidential information or violations of data protection regulations. In addition, the data is considered to have commercial value, and companies do not want to disclose it. To mitigate these risks, all necessary steps must be taken to ensure data security and remove any identifying information. This allows information to be shared without violating client confidentiality, but this procedure requires significant effort from engineers, managers, and lawyers. Though there exist a small number of properly anonymized banking datasets, such as credit card transactions (Padhi et al., 2021) or AlphaBattle (alp, [n. d.]), to the best of our knowledge, there are no publicly available large multimodal temporal datasets for banks.
@@
Thus, this paper introduced the first large-scale multimodal banking dataset to support future research on multimodal techniques for event sequences. In particular, we select several practically essential tasks, such as campaigning (Rosset et al., 2001), i.e., prediction if a client would purchase some of four rather popular products in the next month. Each client is described by sequences of typical bank data: transactions, geo positions where the customer used the bank application, and dialogues with technical support. These data sources highlight the main difficulties in developing multimodal models: asynchronous events in different modalities, various intensities of events, rare/irregular events, and the absence of some modalities for many clients. Based on our dataset, the researchers can study cross-modal connections of sequences from multiple sources at the level of individual events.
@@
Let us discuss the details of the dataset collection procedure. At first, we select a complete sample of clients for two years (2021 and 2022) to cover all seasons. Among all customers who had the opportunity to purchase at least one of four products during 2022, we randomly chose 2,186,230 clients, among which 1M customers are labeled by monthly aggregated purchases of each of four products in each month. For these clients, we collect 947,899,612 financial operations, 1,117,213,760 geo position events, and 5,080,781 dialogues with technical support. Many clients do not have all three modalities simultaneously because they can never make a transaction, call tech support, or leave their geo trace while running the bank application. Next, all the data are properly anonymized to guarantee the confidentiality and privacy of customer information. As a result, it is impossible to recover real clients from our anonymous data. However, we will experimentally show that such anonymization still allows us to extract valuable client information.
We present a comparative analysis of various event sequence datasets alongside the MBD dataset in Table 1. Our dataset is substantially more comprehensive, offering more modalities, events, clients, and downstream tasks than other datasets. MBD incorporates diverse modalities, including bank transactions, geo-locations, and technical support dialogues, providing a richer and more realistic basis for analysis. Additionally, it supports a broader range of downstream tasks, which are detailed in the following sections, allowing for more sophisticated and flexible modeling approaches.
Given the substantial computational resources required for research with our dataset, we have curated a smaller and more balanced subset of MBD (“MBD-small”) to facilitate experimentation under resource-constrained conditions. We included all clients with at least one positive target value from the last month. Then, we also sampled clients with all zeros in their targets at a ratio of 1:10. The balanced subset contains 69474 clients in total. This subset retains the key characteristics of the full dataset while significantly reducing computational overhead.
@@
To demonstrate data transformation from various sources, Fig. 1 briefly shows examples for each modality, and the temporal data structure for a campaigning model is presented. Let us briefly describe each modality in the following subsection. Further details are provided in Appendix A, including the sequence length of event sequences and data samples.
3.2. Modalities
@@
1. Bank transactional data are financial operations carried out between different clients. The main component of our MBD dataset is the client’s transactional history, represented by an event with a timestamp and various attributes of the anonymized counterparty. Clients have 638 transactions on average. Collected over two years (2021 and 2022), the sequence of financial operations can uniquely characterize the client (Babaev et al., 2022), so this data source plays one of the most significant roles in planning and recommendations.
@@
2. Dialog data consists of transcriptions from customer calls to technical support and negotiations between clients and their managers. We incorporate dialogues from key communication channels, including sales and service calls, which account for most interactions with bank customers. It is an extremely important source of information about client needs and problems (Bauman et al., 2024). The audio utterance is fed into a commercial Speech-to-Text library. Personal information, e.g., the client’s name, is detected in the text and masked. To further anonymize the dialogue, we feed its text into a pre-trained NLP model111 https://huggingface.co/ai-forever/ruBert-base and save the resulting embeddings of size 768 in dialogue modality. Only 46% of customers contact support and have records of conversations, and 98% of them have no more than 10 dialogues.
@@
3. Geostream data contains a sequence of geo-coordinates of a client obtained throughout 2022. To anonymize this modality, the coordinates are encoded using geohashes222https://pypi.org/project/pygeohash/, a geocoding system that converts a geographic location into a short string. Each unique geohash corresponds to a region on the Earth’s surface. It is possible to adjust the accuracy and size by removing characters from the end of the code. In our dataset, the coordinates are encoded with a precision of 4, 5, and 6 characters, representing cells of different sizes on the map. As a result, there are 43,999 distinct values of geohash_4, 347,698 numbers of geohash_5, and 2,264,404 most precise locations (geohash_6).
@@
4. Products purchases. Our dataset serves as a valuable resource for analyzing bank customers’ needs and optimizing the campaigning process. This critical task directly impacts the bank’s product volume and overall profitability. High-quality recommendations play a key role in enhancing the customer experience, making campaigning essential for business outcomes and customer satisfaction. Specifically, MBD includes monthly data on the purchases of four distinct banking products throughout 2022, providing a broader temporal scope that captures patterns beyond the pandemic’s peak. We concentrate on the most popular of these products, as internal analysis across various tasks using proprietary data consistently showed that these products provide a robust foundation for model selection. The insights from this data enable the development of models that demonstrate superior performance across a broad spectrum of related tasks. To predict a purchase in a certain month, it is necessary to take events (transactions, geo, dialogues) strictly before the beginning of this month. Therefore, the purchase dataset’s date range is shifted by 1 month, i.e., information is available from February 1, 2022, to January 31, 2023. The campaigning task is a multi-label classification problem, i.e., we store a binary label for each product that indicates whether a customer purchases it in a certain month. The peculiarity of this dataset is its imbalance, which is specific to this type of business task: 81% of clients have no purchases, 15% have one, and the remaining 4% have two or more purchases. A historical overview over 12 months allows us to model the customer behavior dynamic and predict the date of purchase more accurately.
3.3. Data anonymization
@@
Our dataset contains no personal or confidential information whatsoever. Nevertheless, the event sequences are detailed enough that it could be possible to compare individuals from the publicly accessible portion of the dataset with the original proprietary data. To mitigate this risk, noise is introduced to the data, ensuring that such comparisons and identification are impossible. Our bank’s internal security department selected the noise patterns. These patterns are applied locally, preserving the overall structure of the data. The specific noise parameters are not disclosed to prevent potential attacks on the dataset.
@@
All ID fields are hashed with a random salt. All categorical field values are mapped to enumerated indexes. Random noise is added to numerical fields and dates, preserving the hour of the original date, which may cause the local sequence’s shuffle. The dialogue embedding space is divided into regions, which are then shuffled.
In the experimental study for downstream tasks, we demonstrate that the applied anonymization techniques have a minimal impact on the ranking of models trained on our dataset. This indicates that the anonymization process preserves fundamental structural properties while retaining information essential for the downstream task, thereby ensuring the reliability of model evaluation.
4. Benchmark
In this paper, we introduce a benchmark for widely used event sequence datasets, incorporating practically important downstream tasks. This section provides a detailed description of the downstream tasks, baseline methods, and evaluation protocols.
4.1. Datasets and downstream tasks
We implement an out-of-fold validation protocol for each downstream task in every dataset. The client dataset is partitioned into five folds, with four folds used for training and the remaining fold reserved for testing. The training and testing sets are publicly available alongside the dataset, allowing future researchers to compare performance metrics. As each dataset in our benchmark exhibits label imbalance, ROC AUC is the most robust and informative evaluation metric due to its resilience to class imbalances. In real-world business, campaign effectiveness is measured by revenue, but conducting A/B tests for every ML model is impractical. Instead, ROC AUC is a reliable proxy metric for model comparison, with only the top performers advancing to A/B testing. This method was validated through real-world A/B tests and is recognized as the core evaluation metric by leading institutions, including one of the largest global banks. The choice of AUC for campaigning is supported by the reason that ranking models by their ROC curves are similar to comparing their non-response ratio at all possible cutoff points simultaneously (Liu et al., 2012; Rosset et al., 2001). Let us discuss the details of downstream tasks for each dataset in our benchmark.